OpenStudy (anonymous):

sqrt(28) over 9sqrt(16)

OpenStudy (anonymous):

\[\sqrt{28}\over9\sqrt{16}\]

OpenStudy (anonymous):

sqrt 16 = 4 sqrt 28 = 2*sqrt(7) Simplify.

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

what part?

OpenStudy (anonymous):

everything

OpenStudy (anonymous):

you get 9*4 = 36 in the denominator. numerator is 2*sqrt(7). So, you get sqrt(7)/18

OpenStudy (anonymous):

ohhhh :) thankyou!